use factoring to solve the quadratic
x^3-2x^2-x+2=0

Answer by stanbon(75887)

use factoring to solve the quadratic
x^3-2x^2-x+2=0
-----
x^2(x-2)-(x-2) = 0
(x-2)(x^2-1) = 0
---
(x-2)(x-1)(x+1) = 0
----
x = 1 or 2 or -1
===================
Cheers,
Stan H.
===================